Calculate the mass of air in a 100m of floor dimension 10m×12m and height 4m. Take density of air =1.2kg/m^3


1
Expert's answer
2021-01-27T14:28:02-0500

By definition, the mass is:


"m = \rho V"

where "\rho = 1.2kg/m^3" is the density of air, and "V" is the volume of the floor. It is given as the product of the floor dimensions:


"V = 10m\times 12m\times 4m = 480m^3"

Thus, the mass is:


"m = 1.2kg/m^3\times 480m^3 = 576kg"

Answer. 576 kg.